cos 4θ + 4 cos 2θ ≡ 8 cos4 θ − 3.

Answer by mananth(16949)

cos(4θ) =
cos(2*(2θ)) =
2cos^2(2θ) - 1 =
2[ 2cos^2(θ) - 1 ]^2 - 1
2 [ 4cos^4(θ) - 4cos^2(θ) + 1 ] - 1
8cos^4(θ) - 8cos^2(θ) + 1.....................(1)
.............
4 cos 2theta
4(2 cos^2+Theta-1)
8 cos^2+theta-4................................(2)
(1)+(2)= RHS

8cos^4(θ) - 8cos^2(θ) + 1 + 8 cos^2+theta-4= RHS